OpenStudy (anonymous):

The radius of a cylinder is 3.5 ft. The height is 14 ft. Find the surface area and volume of the cylinder to the nearest tenth of a foot. Show your work.

OpenStudy (anonymous):

\[8\times9\pm22\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

SA=\[(2\pi r^2)+(2\pi r*h)\] \[(2\pi 3.5^2)+(2\pi 3.5*14)\] \[24.5\pi +98\pi\] \[=122.5\pi ft^2\] \[V=\pi r^2*h\] \[\pi3.5^2*14\] \[=171.5\pi ft^3\]

OpenStudy (anonymous):

thank you to who